OpenStudy (anonymous):

How to answer: 3/6 = x-3/8 (all of x-3 is divided by 8)? Thanks.

OpenStudy (nubeer):

1/2 = ( x-3)/ 8 8/2 = (x-3) 4 = x-3 x= 7

OpenStudy (ajprincess):

\[\frac{ 3 }{ 6 }=\frac{ x-3 }{ 8 }\] Cross multiply. then u will get 3*8=6(x-3) 24=6*x-6*3 24=6x-18 add 18 both sides 24+18=6x 42=6x divide by 6 both sides 42/6=x x=7

OpenStudy (ajprincess):

Is it clear?

OpenStudy (ajprincess):

@silkroadtomandarin is it clear? or u want further explanation?

OpenStudy (anonymous):

It is very clear. I really appreciate it. I will study these steps.